# ClaraVid: A Holistic Scene Reconstruction Benchmark From Aerial Perspective With Delentropy-Based Complexity Profiling

### v1

@@ 3.4 Delentropic Scene Profile Interpretation
The Delentropic Scene Profile (DSP) quantifies scene complexity through three key statistical descriptors: the mean ($\mu$), standard deviation ($\sigma$), and the shape parameters ($\alpha,\beta$) of the Beta-distributed delentropy values, $H_{\text{del}}\sim\text{Beta}(\alpha,\beta,a,b)$. The mean $\mu$ encapsulates the overall structural complexity of a scene. Low-complexity environments ($\mu<2.5$) predominantly contain large uniform regions, such as clear skies, bodies of water, or uniformly textured surfaces, whereas high-complexity scenes ($\mu>4.75$) are characterized by high frequency textures such as dense vegetation, roof tilings, etc. The standard deviation $\sigma$ quantifies the dispersion of delentropy values across the scene, with higher values indicating greater structural heterogeneity. The shape parameters $\alpha$ and $\beta$ define the skewness of the delentropy distribution: $\alpha<\beta$ signifies a predominance of low-detail imagery, while $\alpha>\beta$ indicates a tendency towards high-complexity views. When $\alpha\approx\beta<1$, the distribution is bimodal, capturing scenes with both extreme simplicity and complexity, whereas $\alpha\approx\beta>1$ suggests a concentration of delentropy values around moderate complexity levels. Together, these descriptors provide a quantitative framework for characterizing aerial scene complexity, facilitating the selection of optimal reconstruction strategies. A visual example is presented in Figure 2.
@@
Discussion In the context of neural scene reconstruction, the DSP provides a lens to interpret different types of predictive uncertainty—epistemic (coverage gaps in training or inherent reconstruction method constraints) and heteroscedastic (unreducible errors). A long tail of low-delentropy images (e.g., large uniform regions such as sky or water) offers minimal geometric cues, intensifying epistemic and model uncertainty. Here, standard strategies like increasing the number of training samples or model size often fail to reduce error; instead, domain-specific solutions—such as incorporating depth priors have the potential to be more effective[52]. Conversely, a long tail of high-delentropy images, images with intricate textures and geometry, (i.e densely tiled roofs, trees, etc) can drive large variations when the model is under-parameterized, and additional viewpoints or higher network capacity to capture subtle details. In both extremes cases of delentropy values, heteroscedastic uncertainty limits how much the error can be reduced solely through oversampling or upscaled architectures. By revealing how complexity is distributed across scenes, the DSP informs targeted resource allocation and methodology choices, that ultimately can improve reconstruction fidelity.
@@ 4 ClaraVid Dataset
In this section we describe ClaraVid, a synthetic dataset designed for semantic and geometric reconstruction tasks. It comprises 16,917 multimodal frames collected across 5 diverse environments—urban, urban high, rural, highway, and natural landscapes— through 8 distinct UAV missions. The dataset is structured to serve as a benchmark for aerial perception tasks, such holistic scene reconstruction and the downstream supporting tasks such as panoptic segmentation, depth estimation, etc.
@@
To ensure comprehensive scene coverage, ClaraVid employs a structured grid-based acquisition strategy, capturing images with an 80–90% overlap to balance redundancy and spatial continuity. Each mission is recorded from 3 distinct viewpoints, simulating a multi-UAV coordination paradigm. Data is acquired at varying altitudes and pitch angles: low (45–55m) 45°, mid (55–65m) 55°, and high (65–75m) 90°, with altitude adjustments dynamically adapting to terrain complexity. Importantly, the dataset does not consist of continuous image sequences; instead, there is an intentional 2–3 second pause between consecutive captures to simulate real-world aerial mapping scenarios, where UAVs acquire individual high-quality frames rather than video streams. ClaraVid spans a total area of 1.8 km2, with an average scene coverage of 0.22 km2, ensuring a rich variety of aerial perspectives.
@@
The dataset provides a diverse set of sensor modalities to support neural reconstruction tasks. High-resolution RGB imagery is captured using a simulated camera at 4032$\times$3024 resolution with a 75 horizontal field of view. To encode geometric priors, ClaraVid includes metric depth maps and scene-level sparse point clouds (color, semantic, instance) sampled at three resolutions (30 m, 100 m, and 200 m), simulating LiDAR scans. For semantic understanding, panoptic segmentation masks provide instance-aware annotations for static structures such as buildings and dynamic entities including vehicles and pedestrians, with a semantic vocabulary of 18 classes. A key feature is the inclusion of dynamic object masks, which isolate moving entities, enabling motion-aware processing—an essential component for tasks requiring temporal coherence, such as NeRFs and dynamic scene reconstruction, to the best of our knowledge, feature unavailable in prior works. These multimodal annotations position ClaraVid as a comprehensive dataset for both static and dynamic aerial perception challenges. We present the delentropic scene profiles of all scenes in Figure 3, for a better understanding of the proposed dataset and provide additional samples in Section 10.
4.1 Development Environment
@@
ClaraVid is developed using Unreal Engine, and leveraging the Carla[53] ecosystem for sensor fidelity and dynamic agent behavior, ensuring diverse yet reproducible traffic patterns. The dataset spans 5 distinct environments. For urban, rural, and highway scenarios, we enhance Carla’s original towns with photorealistic assets, refining textures, lighting, and object placement to elevate scene fidelity beyond default synthetic environments. The natural landscape is procedurally synthesized using real-world elevation [54] and road network priors from OSM[55], ensuring topographic realism, while the high-density urban environment is constructed from professionally designed assets that introduce architectural complexity. Across all scenes, procedural generation enriches foliage and ground cover—introducing dynamic trees, grass, and natural elements—to counteract the overly uniform, planar compositions found in prior synthetic datasets ([12, 13]. Manual refinements ensure spatial coherence and realism, creating environments that balance controlled benchmarking with the visual and structural richness needed for real-world applicability.
5 Experiments
@@
Table 2: Novel view synthesis evaluation on ClaraVid. We compare multiple neural reconstruction methods under varying DSP, characterized by the mean ($\mu$) and shape parameters ($\alpha,\beta$). Higher $\mu$ values strongly correlate with reduced PSNR/SSIM ($\rho\approx-0.68$) and increased LPIPS ($\rho\approx+0.67$), indicating that greater scene complexity degrades reconstruction quality. $\alpha$ shows weak direct correlations, while a higher $\beta$, which skews the distribution toward simpler views, partially offsets performance loss by exhibiting a positive correlation with PSNR/SSIM and a negative correlation with LPIPS ($\rho\approx-0.3$).
@@
Figure 4: Correlation between Delentropy and Reconstruction Score. Delentropy strongly correlates (Pearson) with reconstruction difficulty: higher delentropy corresponds directly to increased perceptual errors (LPIPS) and inversely to texture(PSNR) and geometric metrics(SSIM). The results are computed at image level aggregated across test splits.
Figure 5: Viewpoint Generalization in Neural Reconstruction. Comparison of Nerfacto and 3DGS across RGB, semantics, and depth. Trained on 90°||||H, Nerfacto generalizes better to unseen views (45°||||L, 55°||||M), while 3DGS excels in depth but degrades under shifts. Semantic accuracy follows reconstruction trends, with 3DGS performing well at training views but losing consistency under extrapolation.
5.1 Reconstruction Performance
We evaluate the reconstruction performance of various neural reconstruction models—Nerfacto [57], TensoRF [56], Instant-NGP [36], ZipNeRF [58], and Gaussian Splatting (3DGS) [9]—on the ClaraVid dataset, covering diverse scenarios characterized by distinct $DSP_{s}$. Performance is evaluated using PSNR, SSIM, and LPIPS, with results summarized in Table 2. A clear trend emerges where increased scene complexity, indicated by higher DSP values, leads to a consistent decline in reconstruction fidelity. This is reflected in the strong negative correlation between PSNR/SSIM and DSP mean values ($\rho\approx-0.68$), alongside a positive correlation with LPIPS ($\rho\approx+0.67$), signifying heightened perceptual error in more intricate scenes. While Gaussian Splatting and ZipNeRF demonstrate superior accuracy in structured urban environments, their advantage diminishes in more complex rural and natural landscapes. Interestingly, some exceptions arise. Namely, in natural scenes with substantial structural variation, NeRF-based methods maintain relatively higher PSNR scores, suggesting an ability to capture dominant large-scale structures effectively. However, these same models exhibit weaker perceptual quality, as reflected in their SSIM and LPIPS scores, indicating limitations in preserving fine details. A higher $\beta$ appears to mitigate reconstruction degradation, suggesting that scenes with simpler images are less susceptible to performance drops. Quantitative correlations between delentropy and reconstruction error across datasets further validate delentropy as a robust predictor of reconstruction difficulty (see Figure 4). We recognize the potential limitation of evaluating only synthetic data and explore its real-world applicability, detailed in Section  8.
Table 3: Performance Across Varying Viewpoints . Evaluation on appearance, segmentation, and depth synthesis across various viewpoints: $45^{\circ}|\text{{L}ow}$, $55^{\circ}|\text{{M}id}$, and $90^{\circ}|\text{{H}igh}$. (pitch||||altitude)
@@ 5.1 Reconstruction Performance
5.2 Performance Across Varying Viewpoints
We evaluate the generalization capability of 2 neural reconstruction paradigms—NeRF based NeRF (Nerfacto) and particle-based Gaussian Splatting (3DGS) — in the context of aerial viewpoint variations. This experiment is motivated by real-world UAV-based mapping scenarios, where viewpoint interchangeability is essential (e.g., oblique images intended for perspective mapping may later be repurposed to generate nadir orthophotos). Each method is trained on images from a single altitude angle ($45^{\circ}$, $55^{\circ}$, or $90^{\circ}$) and subsequently tested on unseen viewpoints to assess robustness under realistic operational conditions. The quantitative results, detailed in Table 3, reveal distinct performance profiles across novel view synthesis, semantic segmentation (mIoU), and depth estimation (RMSE, AbsRel, and $\delta^{1.05}$). Nerfacto exhibits greater robustness to viewpoint variations, maintaining relatively consistent reconstruction quality across novel views, albeit with longer training times. Conversely, 3DGS achieves higher fidelity on familiar views and converges faster but degrades significantly under moderate viewpoint shifts, limiting its suitability for applications requiring broad viewpoint coverage. Semantic segmentation accuracy (mIoU) closely follows reconstruction quality trends, with 3DGS performing better at training viewpoints but losing accuracy under extrapolation, suggesting a strong dependence on appearance consistency. Depth estimation results, however, favor 3DGS, which consistently outperforms Nerfacto, indicating superior geometric reconstruction. We used a stricter depth accuracy threshold of $\delta<1.05$, as even minor inaccuracies can propagate into large errors at great distances ($>100m$). For visual comparison see Figure 5.
@@ 5.2 Performance Across Varying Viewpoints
Syndrone[12] Syndrone 0.051 86.36
@@ 5.3 Semantic Segmentation as downstream task
We assess ClaraVid’s effectiveness in downstream tasks by quantifying the synt-to-real domain gap in semantic segmentation. A segmentation model (see Section 9 for details) is trained on real [24, 22, 30] and synthetic datasets, with results shown in Table 4. Despite containing 4.5$\times$ fewer samples than Syndrone [12], ClaraVid achieves a smaller domain gap, demonstrating superior scene complexity and image fidelity. Models trained on ClaraVid consistently outperform those trained on Syndrone across all real datasets, achieving higher mIoU and lower prediction entropy. However, a residual semantic gap remains, with ClaraVid exhibiting an average gap of 17.82%, compared to 22.34% for Syndrone.
@@ 6 Conclusions
In this work, we introduced ClaraVid, a synthetic aerial dataset designed to bridge the gap in holistic scene reconstruction tailored for UAV-based scene understanding. We further proposed the Delentropic Scene Profile, a novel complexity metric rooted in differential entropy, which quantifies scene difficulty and exhibits strong correlations with reconstruction performance. Our benchmarking framework systematically evaluates neural reconstruction techniques, revealing how DSP serves as a reliable predictor of reconstruction accuracy, informing method selection and optimization strategies. Through extensive experimentation, we demonstrate that ClaraVid not only surpasses existing synthetic datasets in realism and annotation quality but also generalizes better to real-world UAV imagery, challenging the conventional notion that dataset size alone dictates generalization. By releasing ClaraVid, we aim to advance UAV-based perception and improve benchmarking for semantic and geometric reconstruction.
@@
7 Delentropic Scene Profile Real Data
To evaluate the effectiveness of the Delentropic Scene Profile (DSP) in real-world environments beyond synthetic datasets, we follow the experimental framework of CityGaussian [45] and assess DSP on urban mapping datasets, including UrbanScene3D [11] and Mill19 [46]. These datasets feature diverse urban landscapes with varying degrees of structural complexity, allowing us to analyze the correlation between delentropy and reconstruction performance under real-world conditions. By computing the delentropy profile for each scene and aggregating the results, we establish a direct comparison with the experimental numbers reported in these studies. As shown in  Figure 6, the scenes exhibit moderate complexity, with the residential scene in particular displaying a DSP profile comparable to the Urban High category in the Claravid dataset. Additionally, we examine large-scale NeRF-based methods, such as Mega-NeRF [46], Switch-NeRF [59], CityGaussian [45], Gaussian Splatting [7], and GN-NeRF [60]. Our findings, present in Figure 7, reveal that higher DSP values are consistently associated with increased reconstruction errors, reinforcing delentropy’s predictive capability in real-world scenarios. These results establish DSP as a robust metric for quantifying scene complexity and estimating reconstruction difficulty.

### v2

@@ 3.3 Delentropic Scene Profile Interpretation
DSP quantifies scene complexity through four key analytical descriptors: the mean $\mu$, standard deviation $\sigma$, and the shape parameters $\alpha,\beta$ of the Beta-distributed delentropy values, $H_{\text{del}}\sim\text{Beta}(\alpha,\beta,a,b)$. The mean $\mu$ encapsulates the overall structural complexity of a scene. Low-complexity environments ($\mu<2.5$) predominantly contain large uniform regions, such as clear skies, bodies of water, or uniformly textured surfaces, whereas high-complexity scenes ($\mu>4.75$) are characterized by high frequency textures such as dense vegetation, roof tilings, etc. The standard deviation $\sigma$ quantifies the dispersion of delentropy values across the scene, with higher values indicating greater structural heterogeneity. The shape parameters $\alpha$ and $\beta$ define the skewness of the delentropy distribution: $\alpha<\beta$ signifies a predominance of low-detail imagery, while $\alpha>\beta$ indicates a tendency towards high-complexity views. When $\alpha\approx\beta<1$, the distribution is bimodal, capturing scenes with both extreme simplicity and complexity, whereas $\alpha\approx\beta>1$ suggests a concentration of delentropy values around moderate complexity levels. Together, these descriptors provide a quantitative framework for characterizing aerial scene complexity, facilitating the selection of optimal reconstruction strategies.
@@
Discussion. In the context of neural scene reconstruction, the DSP provides a lens to interpret different types of predictive uncertainty—epistemic (coverage gaps in training or inherent reconstruction method constraints) and heteroscedastic (irreducible errors). A long tail of low-delentropy images (e.g., large uniform regions such as sky or water) offers minimal geometric cues, intensifying epistemic and model uncertainty. Here, standard strategies like increasing the number of training samples or model size often fail to reduce error; instead, domain-specific solutions—such as incorporating depth priors–have the potential to be more effective[57]. Conversely, a long tail of high-delentropy images, with intricate textures and geometry, (e.g., densely tiled roofs, trees, etc) can drive large variations when the model is under-parameterized, and additional viewpoints or higher network capacity to capture subtle details. In both extremes cases of delentropy values, heteroscedastic uncertainty limits how much the error can be reduced solely through oversampling or upscaled architectures. By revealing how complexity is distributed across scenes, the DSP informs targeted resource allocation and methodology choices, that ultimately can improve reconstruction fidelity.
Limitations. Delentropy inherently relies on image gradients, rendering it sensitive to factors such as blur, resolution variations, or imaging artifacts. In such cases, it becomes unclear whether reconstruction should aim to reproduce the degraded observations or an idealized, high-fidelity version of the scene. Additionally, DSP is formulated under structured acquisition settings, assuming bounded scenes and regular grid-based sampling. When these assumptions break down–for instance, in unbounded environments with significant peripheral content or in scenes requiring spatial contraction–the predictive power of DSP with respect to reconstruction (evaluated via novel view synthesis) quality may deteriorate.
Please consult the Supplementary 8 for more information regarding design decisions of DSP.
@@ 4 ClaraVid Dataset
In this section we describe ClaraVid, a synthetic dataset designed for semantic and geometric reconstruction tasks. It comprises 16,917 multimodal frames collected across 5 diverse environments—urban, urban high, rural, highway, and natural landscapes— through 8 distinct UAV missions. ClaraVid spans scenes of varying structural complexity, from uniform environments like Urban or Highway to highly detailed ones such as Rural and Nature. This diversity supports controlled benchmarking across different levels of reconstruction difficulty. We present the delentropic scene profiles in Figure 3.
@@
Data collection. To ensure comprehensive scene coverage, ClaraVid employs a structured grid-based acquisition strategy, capturing images with an 80–90% overlap to balance redundancy and spatial continuity. Each mission is recorded from 3 distinct viewpoints, simulating a multi-UAV coordination paradigm. Data is acquired at varying altitudes and pitch angles: low (45–55m) 45°, mid (55–65m) 55°, and high (65–75m) 90°, with altitude selected based on terrain complexity. Importantly, frames are captured at 2–3 second intervals, reflecting real-world UAV mapping protocols that prioritize discrete high-quality images over continuous video. ClaraVid spans a total area of 1.8 km2, with an average scene coverage of 0.22 km2.
@@
Data modalities. The dataset provides a diverse set of sensor modalities to support neural reconstruction tasks. High-resolution RGB imagery is captured using a resolution of 4032$\times$3024 with a 75 horizontal FoV. To encode geometric priors, ClaraVid includes metric depth maps and scene-level sparse point clouds (color, semantic, instance) sampled at three densities (30, 100 and 200 cm), simulating LiDAR scans. For semantic understanding, panoptic segmentation masks provide instance-aware annotations for static structures such as buildings and dynamic entities including vehicles and pedestrians, with a semantic vocabulary of 18 classes. A key feature is the inclusion of dynamic object masks, which isolate moving entities, an essential component for tasks requiring temporal coherence. To the best of our knowledge, this feature is unavailable in prior works. These multimodal annotations position ClaraVid as a comprehensive dataset for aerial perception challenges.
@@
Development Environment. ClaraVid is developed using Unreal Engine 4, and leveraging the Carla[58] ecosystem for sensors and dynamic agent behavior, ensuring diverse traffic patterns. The dataset spans 5 distinct environments. For urban, rural, and highway scenarios, we enhance Carla’s original towns with photorealistic assets, refining textures, lighting, and object placement to elevate scene fidelity beyond default synthetic environments. The natural landscape is created using real-world elevation [59] and road network priors from OSM[60], ensuring topographic realism, while the high-density urban environment is constructed from professionally designed assets that introduce architectural complexity. Across all scenes, procedural generation enriches foliage and ground cover–introducing trees, grass, and natural elements–to counteract the overly uniform, planar compositions found in prior synthetic datasets ([12, 13]. Manual refinements ensure spatial coherence and realism, creating environments that balance controlled benchmarking with the visual and structural richness needed for real-world applicability.
Additional information about the dataset can be found in Supplementary 11.
5 Experimental Evaluation
@@
Figure 4: Correlation between Complexity Metrics and Reconstruction Error. Delentropy consistently achieves stronger Pearson correlations and tighter predictive intervals with reconstruction metrics compared to GLCM texture entropy and Shannon pixel entropy. Higher delentropy corresponds directly to increased perceptual errors (LPIPS) and inversely to texture(PSNR) and geometric metrics(SSIM). The results are computed at image level aggregated across test splits.
5.1 Reconstruction Performance
Table 2: Novel view synthesis evaluation on ClaraVid. We benchmark several neural reconstruction methods across scenes with different Delentropic Scene Profiles. A higher $\mu$ consistently correlates with lower PSNR/SSIM and higher LPIPS, confirming that increased scene complexity degrades reconstruction quality. While $\alpha$ has limited predictive value, higher $\beta$—indicating skew toward simpler views—partially mitigates performance loss.
@@ 5.1 Reconstruction Performance
We evaluate the reconstruction performance of various neural reconstruction models–Nerfacto [62], TensoRF [61], Instant-NGP [40], ZipNeRF [63], and Gaussian Splatting [7]–on the ClaraVid dataset, covering diverse scenarios characterized by distinct DSP. Performance is evaluated using PSNR[64], SSIM[65], and LPIPS[66], with results summarized in Table 2. Full experimental setup can be found in Supplementary  10.1. A clear trend emerges where increased scene complexity, indicated by higher DSP values, leads to a consistent decline in reconstruction fidelity. This is reflected in the strong negative correlation between PSNR/SSIM and DSP mean values, alongside a positive Pearson correlation with LPIPS. A higher $\beta$ appears to mitigate reconstruction degradation, suggesting that scenes with simpler images are less susceptible to performance drops. Among all methods, Gaussian Splatting and ZipNeRF deliver the most robust performance across settings.
5.2 DSP as complexity proxy
To evaluate delentropy as a proxy of reconstruction difficulty, we analyze its correlation with reconstruction error across two models–Gaussian Splatting and Nerfacto–on the ClaraVid dataset. Following the setup from the previous subsection, DSP is computed on grayscale frames prior to training, excluding masked dynamic regions. As shown in Figure 4, delentropy achieves consistently higher correlation with reconstruction metrics than classical complexity measures such as Shannon entropy[51] and GLCM texture entropy[52]. It correlates negatively with PSNR and SSIM, and positively with LPIPS, indicating higher reconstruction errors in structurally complex regions. This trend is consistent across both models, suggesting that delentropy captures structural cues relevant to reconstruction regardless of model class. Moreover, delentropy yields tighter predictive intervals, validating its reliability as a proxy for reconstruction complexity and motivating its use in the Delentropic Scene Profile.
5.3 Performance Across Varying Viewpoints
Figure 5: Viewpoint Generalization in Neural Reconstruction. Comparison of Nerfacto and 3DGS across RGB, semantics, and depth. Trained on 90°|||H, Nerfacto generalizes better to unseen views (45°|||L, 55°|||M), while 3DGS excels in depth but degrades under shifts. Semantic accuracy follows reconstruction trends, with 3DGS performing well at training views but losing consistency under extrapolation.
Table 3: Performance Across Varying Viewpoints . Evaluation on appearance, segmentation, and depth synthesis across various viewpoints: $45^{\circ}|\text{{L}ow}$, $55^{\circ}|\text{{M}id}$, and $90^{\circ}|\text{{H}igh}$. (pitch|||altitude)
@@ 5.3 Performance Across Varying Viewpoints
[…]
@@
Syndrone Syndrone 0.051 86.36
Dronescapes Dronescapes 0.153 52.74
@@ 5.4 Semantic Segmentation as Downstream Task
We assess ClaraVid’s effectiveness in downstream tasks by quantifying the synt-to-real domain gap in semantic segmentation. A DPT-based model is trained on real [25, 22, 24, 26] and synthetic datasets [12], with results shown in Table 4. Despite containing 4.5$\times$ fewer samples than Syndrone [12], ClaraVid achieves a smaller domain gap, demonstrating superior scene complexity and image fidelity. Models trained on ClaraVid consistently outperform those trained on Syndrone across all real datasets, achieving higher mIoU and lower prediction entropy. However, a residual semantic gap remains, with ClaraVid exhibiting an average gap of 17.82%, compared to 22.34% for Syndrone.
@@ 6 Conclusions
In this work, we introduced ClaraVid, a synthetic aerial dataset designed to bridge the gap in holistic scene reconstruction tailored for UAV-based scene understanding. We further proposed the Delentropic Scene Profile, a novel complexity metric rooted in differential entropy, which quantifies scene difficulty and exhibits strong correlations with reconstruction performance. Our benchmarking framework systematically evaluates neural reconstruction techniques, revealing how DSP can be a reliable predictor of reconstruction accuracy. Through extensive experimentation, we demonstrate that ClaraVid not only surpasses existing synthetic datasets in realism and annotation quality but also generalizes better to real-world UAV imagery, challenging the conventional notion that dataset size alone dictates generalization. By publicly releasing ClaraVid, we aim to advance UAV-based perception and improve benchmarking for semantic and geometric reconstruction.
7 Acknowledgment
This work was supported by the project ”Romanian Hub for Artificial Intelligence-HRIA”, Smart Growth, Digitization and Financial Instruments Program, MySMIS no. 334906.
@@ 7 Acknowledgment
8 More on Delentropic Scene Profile
This section expands on the design and interpretation of the Delentropic Scene Profile. We justify the use of the Beta distribution and clarify the conditions required for reliable cross-scene comparison. These clarifications aim to guide appropriate usage of DSP as a diagnostic tool in structured 3D scene understanding.
8.1 Motivation for Beta Distribution
The choice of distributional model for $\text{DSP}_{S}$ was guided by the analysis of complexity profiles of diverse datasets—ClaraVid, Skydrone[12], Skyscapes[24], UAVid[25], 3D Matrix City [10], Mill19 [50], and UrbanScene3D [11]—observing varied distributional forms contingent on the reconstruction context. A Gaussian distribution, while apt for uniformly complex scenes, fails to capture skewed or long-tailed distributions prevalent in heterogeneous environments (for example where we have a body of water or sky regions). Alternatives such as the Pareto or Gamma distributions accommodate skewness but lack flexibility for U-shaped profiles, which emerge in scenarios with pronounced bimodal complexity (e.g., a field with occasional high-detail trees). The Beta distribution, however, adeptly models this spectrum—uniform, skewed, and U-shaped—owing to its shape parameters $\alpha$ and $\beta$, requiring only four parameters (including bounds $a$ and $b$) for a parsimonious yet expressive fit. This simplicity affords intuitive interpretation: $\alpha$ and $\beta$ directly govern the density’s form, while $a$ and $b$ anchor it to the empirical delentropy range. Multimodal distributions were eschewed, as their complexity—necessitating mixture models with additional parameters—hampers interpretability, risks overfitting, particularly given the finite sample sizes of $\{H_{\text{del},k}\}$. The Beta distribution emerges as the best choice, balancing flexibility, robustness, and theoretical clarity.
8.2 Comparability Across Scenes
The Delentropic Scene Profile provides a statistical summary of structural complexity and is designed to support comparisons of scenes in terms of their representational difficulty. While delentropy provides an absolute measure of complexity, its reliability for cross-scene comparison hinges on controlled acquisition conditions. Image resolution plays a critical role: small variations (e.g., $\approx$10%) tend to preserve the stability of the DSP, whereas more aggressive change of scales (e.g., $\geq$25%) can significantly alter high-frequency content, shifting the delentropy distribution (e.g., roof tiles may become indistinct when downsampled). The spatial extent of the covered scene is similarly influential. We have observed that DSP remains stable even under 4× scaling in area, but larger deviations or differences in semantic composition introduce shifts in the aggregated profile due to increased structural heterogeneity. Furthermore, the image collection policy–particularly the trajectory and sampling layout–can induce biases in the captured content, making structurally similar scenes appear dissimilar under mismatched acquisition patterns. These observations emphasize that to fully leverage DSP as a comparative tool, one should ensure alignment in resolution, spatial coverage, and collection policy—allowing differences in DSP to reflect genuine variations in scene complexity rather than artefacts of sampling.
8.3 Implementation Details
The delentropy of an image is obtained by first applying Gaussian blur to the input image using a spatial kernel of $\times 3$ 3 pixels with a standard deviation of 1.01.01.0, reducing high-frequency noise and mitigating sensitivity to minor textural variations. Subsequently, spatial gradients along horizontal and vertical directions are computed via Sobel filters with kernel size $\times 3$ 3. The resulting gradient field is quantized into a two-dimensional histogram, the deldensity, employing 256256256 bins per axis. This choice is driven by practical considerations to avoid histogram saturation, an artifact typically encountered when considering both positive and negative gradient values. Finally, the normalized deldensity is used as the joint probability distribution, from which delentropy is calculated.
9 Supplementary Evaluation for DSP
In this section, we evaluate the DSP on real-world datasets to assess its robustness under varied capture policies and reconstruction conditions.
9.1 UAVid
To evaluate the transferability and generalization of the DSP to real-world UAV imagery, we assess its correlation with reconstruction accuracy on the UAVid dataset. Unlike ClaraVid, UAVid features continuous image sequences captured under linear, L-shaped, and U-shaped trajectories, with no explicit grid-based coverage. We select six scenes (seq_13, seq_15, seq_29, seq_31, seq_36, seq_38), prioritizing static intervals. Full experimental setup provided in Section  10.4.
Despite a reduction in correlation between DSP and reconstruction metrics, the results–summarized in Table 5–remain consistent with prior trends observed in synthetic settings as presented in Figure 6. Notable degradations are attributed to real-world artifacts such as rolling shutter, motion blur, and the absence of structured spatial coverage, which collectively weaken the delentropy signal. We identify two recurring failure modes: (i) blurred input frames that yield reconstructions of unexpectedly high fidelity, and (ii) insufficient coverage in low-altitude flyovers (e.g., single-pass views) that lead to poor reconstructions despite an average measured delentropy. Visual examples provided in Figure 7.
Table 5: DSP evaluation on UAVid. Higher DSP $\mu$ correlates with lower PSNR, SSIM and higher LPIPS.
Scene DSP Nerfacto[62] Gaussian Splatting [7]
$\mu$ $\alpha$ $\beta$ PSNR ↑ SSIM ↑ LPIPS ↓ PSNR ↑ SSIM ↑ LPIPS ↓
seq 38 3.984 6.147 2.216 22.83 0.637 0.536 25.18 0.857 0.234
seq 36 3.990 1.264 0.812 22.80 0.543 0.493 25.10 0.818 0.215
seq 29 4.017 3.187 1.702 21.93 0.525 0.512 24.07 0.770 0.238
seq 13 4.018 6.911 2.159 21.41 0.497 0.578 23.70 0.730 0.326
seq 31 4.033 1.470 1.659 21.75 0.463 0.475 23.93 0.718 0.191
seq 15 4.048 9.008 2.300 21.20 0.487 0.547 22.98 0.709 0.350
9.2 DSP Analysis on Large Reconstructions
To further assess the generalization of DSP in real-world conditions, we analyze datasets used in recent large-scale urban reconstruction works, following the experimental setup of CityGaussian [49]. Specifically, we compute DSP values on the training and evaluation splits of UrbanScene3D [11] and Mill19 [50], which capture diverse urban geometries under unconstrained conditions. These datasets include varying levels of structural complexity, from sparse residential layouts to dense industrial sites. We align our computed DSP statistics–presented in Figure 8–with published reconstruction scores reported in CityGaussian, enabling a post-hoc correlation analysis without re-running the experiments.
As illustrated in Figure 9, we observe a general trend whereby scenes with higher DSP values correspond to lower reconstruction quality, particularly in PSNR. For example, the Mill19 industrial scene exhibits elevated delentropy ($\mu$ >3.8>3.8> 3.8) and lower PSNR across multiple reconstruction methods, comparable to Urban High scenes in ClaraVid. This negative correlation persists across architectures–Mega-NeRF [50], CityGaussian [49], GN-NeRF [67], Switch-NeRF [68], and Gaussian Splatting [7]–suggesting that DSP offers a dataset-agnostic signal of structural difficulty. SSIM and LPIPS exhibited more variability, and their correlation with DSP was less pronounced than that of PSNR. The consistency of trends across methods and environments reinforces DSP’s utility as a lightweight proxy for estimating reconstruction difficulty in real-world benchmarks.
Figure 6: Correlation between complexity metrics and reconstruction quality on UAVid. Delentropy (blue) exhibits the strongest and most consistent correlation with PSNR, SSIM, and LPIPS across both evaluated models, outperforming GLCM texture entropy (green) and Shannon pixel entropy (red). Metrics are computed per image and aggregated over the test split.
Figure 7: Failure cases. Top: Despite low delentropy and detailed Gaussian Splatting reconstruction, large errors arise from poor camera registration in the blurred input. Bottom: Incomplete scene coverage degrades performance; additionally, the ground truth image exhibits mild rolling shutter distortion.